OpenStudy (anonymous):

The values (7.1, 4) and (x, 11) are from an inverse variation. Find the missing value and round to the nearest hundredth.

OpenStudy (cwrw238):

in verse variation so y = k/x 4= k / 7.1 giving k = 28.4 so y = 28.4 / x for y = 11 11= 28.4 / x x = 28.4 / 1 = 2.58

OpenStudy (anonymous):

Ok so x is 28.4?

OpenStudy (cwrw238):

no x = 2.58

OpenStudy (anonymous):

@cwrw238 When your done could you help me with my recent question

OpenStudy (anonymous):

Ok I Got Confused with the last part

OpenStudy (cwrw238):

yea - ill take a look lawrence

OpenStudy (anonymous):

ty